OpenStudy (anonymous):

If f(x) = 4x – 7 and g(x) = 2x + 4, evaluate f(x) + g(x) for x = –3. A. –52 B. –21 C. –22 D. –15

Nnesha (nnesha):

first step is same like last one add both function and then replace x by -3

OpenStudy (anonymous):

f(x)+g(x)= ( 4x – 7)+( 2x + 4) =6x-3 now substituting x=3 6(3)-3=15

OpenStudy (anonymous):

sorry subtituting x=-3 =6(-3)-3= -18-3=-21

OpenStudy (anonymous):

thanks!